# Unleashing the Potential of Tracklets for Unsupervised Video Person Re-Identification

### v1

@@ IV-B Implementation Details
AGRL [15] 81.9 89.5 96.6 - 95.4 97.0 99.3 -
VRSTC [18] Labels 82.3 88.5 96.5 97.4 93.5 95.0 99.1 99.4
STGCN [14] 83.7 89.9 96.4 - 95.7 97.3 99.3 -
DGM+IDE [22] 16.9 36.8 54.0 - 33.6 42.4 57.9 68.3
Stepwise [23] 19.6 41.2 55.5 - 46.7 56.2 70.3 79.2
Race [24] Label 24.5 43.2 57.1 62.1 - - - -
EUG [21] 42.4 62.6 74.9 - 63.2 72.7 84.1 -
ProLearn [19] 42.6 62.8 75.2 80.4 63.3 72.9 84.3 88.3
DAL [26] Cameras 23.0 49.3 65.9 72.2 - - - -
UTAL [25] 35.2 49.9 66.4 - - - - -
UAAL [20] 60.1 73.2 86.3 89.5 87.0 89.7 97.0 98.6
uPMnet [17] - - - - 76.9 83.6 93.1 -
OIM [60] None 13.5 33.7 48.1 54.8 43.8 51.1 70.5 76.2
TSSL [34] 30.5 56.3 - - 64.6 73.9 - -
DBC [36] 31.7 58.5 70.1 73.5 67.4 75.6 88.5 91.0
BUC [33] 38.0 61.1 75.1 80.0 61.9 69.2 81.1 85.8
NHAC [35] 40.1 61.8 75.3 79.9 76.0 82.8 92.7 95.6
SSL [37] 43.6 62.8 77.2 80.1 69.3 76.4 88.7 91.0
SRC [38] 40.5 62.7 76.1 80.0 76.5 83.0 93.3 95.0
RPE [39] 40.4 63.3 75.4 80.6 71.5 77.8 89.3 91.7
SSR-C (Ours) None 78.7 85.6 93.0 95.1 95.3 95.9 99.2 99.6
IV-C Comparison with the State-of-the-Art
@@
In Tab. I, we present a comparative analysis of our proposed SSR-C against state-of-the-art methods on two large-scale video-based person Re-ID datasets. Our proposed SSR-C surpasses all state-of-the-art unsupervised methods, which overlook not only the unexpected noises and variations within tracklets but also abundant useful identity information in tracklets, by a large margin without whistles and bells. For instance, SSR-C surpasses SRC [38], which resamples and re-weights the hard frames in tracklets to improve clustering ability, with 18.8%/12.9% mAP/Rank-1 on Duke-V, and surpasses SSL [37], which utilizes re-assigned soft label distribution to learn from similar images with smooth constraint, with 35.1%/22.8% mAP/Rank-1 on MARS. Relatively speaking, one-shot learning based and tracklet association based methods usually achieve better performance as shown in Tab. I, partly attributed to their utilization of a few identity annotations or camera metadata for guidance. However, our proposed SSR-C still shows great superiority over them in performance, without relying on any auxiliary information. For example, on MARS, our proposed method significantly outperforms the state-of-the-art one-shot learning based method ProLearn [19], which uses a progressive sampling strategy during the training process, and improves mAP by 36.1%. SSR-C also shows great superiority to the advanced tracklet association based method UAAL [20], which uses the association ranking method to discover positive and negative tracklets pairs for triplet loss training, with an improvement of 18.6% in mAP. It is worth noting that, without expensive annotations or cumbersome modules, our method is competitive with advanced supervised methods [15, 18, 14] on both datasets. The results demonstrate that SSR-C is capable of generating reliable pseudo labels to achieve effective feature learning and improve the discriminative ability of the ReID model.
More comparison with image-based methods. In Tab. II, we also compare with state-of-the-art image-based unsupervised person Re-ID methods [30, 28, 29] by replicating their results on the video-based MARS and Duke-V datasets. Note that we adopt tracklet partition for all these image-based methods on Duke-V, otherwise they struggle to converge as observed in Tab. III (see detailed discussion in Sec. IV-D). Among them, SPCL [28] adopts hybrid memory and improves clustering with reliability criterion. Differently, SSR-C promotes unsupervised video person Re-ID with reliable self-supervised refined clustering and our proposed CSC loss. SSR-C surpasses SPCL significantly on MARS and Duke-V. SSR-C also beats the state-of-the-art unsupervised method PPLR [29], which utilizes both global and local features, with the improvement of 25.1% and 9.2% mAP on MARS and Duke-V, respectively. Moreover, SSR-C surpasses ICE [30], which utilizes additional camera information, by a large margin on both benchmarks.
TABLE II: Comparison with image-based methods on the MARS and Duke-V datasets. “Cameras” denotes requiring camera labels. Our proposed tracklet partition is adopted for comparison methods on Duke-V to ensure convergence.
Methods Extra Info. MARS Duke-V
mAP Rank-1 mAP Rank-1
ICE [30] Cameras 72.3 81.0 86.6 88.2
SPCL [28] None 44.7 59.5 77.0 79.5
PPLR [29] None 53.6 65.9 86.1 86.2
SSR-C (Ours) None 78.7 85.6 95.3 95.9
@@
TABLE III: Ablation studies on the MARS and Duke-V datasets. “SSR” denotes our proposed self-supervised refined clustering, which first generates reliable sub-clusters and then merges them self-supervised. “R./D.R.” and “PM” refer to merging reachable/directly reachable sub-clusters throughout training and the progressive merging strategy, respectively.
@@
IV-D Ablation Studies
@@
In this section, we first investigate the effectiveness of each component in our proposed SSR-C framework through detailed ablation experiments in Tabs. III, IV, V and VI. Then, to help understand our proposed method and demonstrate its superiority, we conduct in-depth analyses of different values of hyper-parameters and their impacts in Figs. 4 and 5, and compare the clustering statistics between the baseline and our method in Fig. 6.
@@
Effectiveness of NFTP. Taking the noises and unexpected variations within tracklets into account, our proposed NFTP module enables reliable clustering. As shown in Tab. III, Method 2 with NFTP brings 10.4% mAP improvement over Method 1 (Baseline) on MARS, and a more remarkable performance improvement is observed on Duke-V. However, directly applying InfoNCE loss based on the original tracklets fails to converge. We argue that tracklets in Duke-V are much longer than MARS (167.6 v.s. 59.5 frames per tracklet on average), which implies more severe variations in tracklets. If we simply average frames within each tracklet without NFTP, the tracklet features with bias would result in inaccurate pseudo labels, and further cause the degradation of the optimization. In contrast, partitioning each tracklet yields sub-tracklets with relatively mild feature changes, which facilitates the feature-based clustering algorithm to obtain accurate pseudo labels. Given that applying NFTP may also increase the amount of training data, we further investigate its influence in Tab. V. The results demonstrate that simply increasing the sampling number of frames has little impact on performance, further demonstrating the effectiveness of NFTP.
@@
Effectiveness of CSC loss and self-supervised refined clustering. Based on the abundant pseudo labels generated with our self-supervised refined clustering, we design CSC loss for effective optimization with dynamically mined positive samples. As shown in the shaded area in Tab. III, compared with Method 2, incorporating either R. or D.R. sub-cluster merging strategy and training with InfoNCE loss show inferior results on MARS. The reason is that although our proposed self-supervised refined clustering produces reliable optimization guidance, InfoNCE still optimizes the model in the coarse tracklet granularity after violent sub-cluster merging. Besides, fusing sub-clusters can inevitably involve negative identities, especially when the capability of the model is weak at the early training stage. However, as the results of Methods 5 and 6 show, our proposed CSC loss can take full advantage of the abundant supervision $\hat{y}$ and $\tilde{y}$ from self-supervised refined clustering. The two complement and promote each other, enhancing the discriminative ability of the model significantly.
@@
Effectiveness of progressive merging strategy. Compared with Methods 5 and 6, Method 7 (Ours) yields an additional improvement on both datasets. The results demonstrate the effectiveness of our progressive merging strategy in adapting the optimization to the feature extraction capability of the model at different training stages.
@@
TABLE IV: Influence of noise filtering and tracklet partition in our proposed NFTP module.
@@
Baseline $+$ NF 71.4 80.7 7.3 9.5
Baseline $+$ TP 72.5 79.8 87.4 89.0
Baseline $+$ NFTP 74.3 82.2 91.7 93.6
SSR-C -- NF 73.8 81.9 93.8 94.6
@@
More ablation studies of noise filtering and tracklet partition. We provide more detailed ablation results of our proposed NFTP module, which contains two key designs of noise filtering and tracklet partition. (1) Noisy filtering is to handle harmful frames as depicted in Fig. 1 (a). In unsupervised learning, noisy frames in tracklets can lead to inaccurate clustering and wrong pseudo labels, thereby hurting model learning. More intuitively, adopting noisy filtering to the baseline in Tab. IV shows its effectiveness. (2) Tracklet partition is designed to alleviate unexpected variations within tracklets as depicted in Fig. 1 (b). Moreover, it helps explore locality within tracklets and enables our progressive self-supervised refined clustering for reliable pseudo labels. We only adopt tracklet partition to the baseline and SSR-C in Tab. IV, and the results show its non-negligible necessity. (3) To promote learning and avoid the potential loss of complementary features within tracklets, we propose a progressive merging strategy to merge sub-tracklets of the same identity effectively. We have shown its effectiveness in Tab. III.
TABLE V: The impact of the different numbers of sampled frames per tracklet (# Frames) on the Duke-V dataset.
Methods # Frames mAP Rank-1 Rank-5
Baseline 8 3.2 4.0 7.1
16 8.5 11.5 19.2
Baseline $+$ TP 4 87.2 88.5 97.2
Baseline $+$ NFTP 4 92.4 93.7 98.4
SSR-C 4 93.5 94.3 99.0
8 95.3 95.9 99.2
16 94.9 95.3 99.3
@@
Influence of the amount of sampled frames. In Tabs. III and IV, we notice a remarkable performance improvement on the Duke-V dataset after applying our proposed NFTP module (especially its tracklet partition design), while the baseline based on the original tracklets fails to converge. Given that applying NFTP may also increase the amount of training data, we further investigate the influence of the number of sampled frames and NFTP on the Duke-V dataset in Tab. V. Firstly, we try different numbers of sampled frames for each tracklet using the baseline method without applying NFTP. The results show that simply increasing the number of sampled frames for each tracklet is useless for enabling model convergence. However, when incorporating tracklet partition, which alleviates the negative impact of unexpected variations within tracklets, a huge performance improvement is observed. Note that on Duke-V, sampling 16 frames per tracklet (Line 2) is roughly equal to the data volume of sampling 4 frames per tracklet after partition (Line 3), but the performance of the former is still unsatisfactory. The results indicate that simply increasing the number of sampled frames almost has no effect on optimization, emphasizing the indispensable role of the tracklet partition design. It alleviates the tracklet feature bias by dividing the tracklets into sub-tracklets and helps improve clustering results for more accurate pseudo labels. What’s more, our proposed NFTP module can better promote learning with noisy frames being filtered, as shown in Line 4 of Tab. V.
We also explore the influence of the different numbers of sampled frames after incorporating NFTP in our proposed SSR-C framework in Tab. V. Note that after applying NFTP, we sample frames in the sub-tracklet level, rather than the original tracklet. Considering the training costs and performance, we sample 8 frames for each sub-tracklet in our experiments. It is worth mentioning that with various numbers of sampled frames, we all achieve impressive results, greatly surpassing existing state-of-the-art methods. The results demonstrate the effectiveness of our proposed SSR-C framework.
Figure 4: Influence of different values of (a) $l$, (b) $K$, (c) $\lambda$. The ablation experiments are performed on the Duke-V dataset and the optimal hyper-parameter values are directly applied to MARS without further tuning.

### v2

@@ IV-B Evaluation Metrics
VRSTC [20] 2019 82.3 88.5 96.5 97.4 93.5 95.0 99.1 99.4
AGRL [17] 2020 81.9 89.5 96.6 - 95.4 97.0 99.3 -
STGCN [16] 2020 Labels 83.7 89.9 96.4 - 95.7 97.3 99.3 -
AGNet [62] 2022 84.3 89.4 96.7 - 96.0 96.3 99.4 -
GPNet [63] 2023 85.1 90.2 96.8 - 96.1 96.3 99.6 -
DGM+IDE [24] 2017 Label 16.9 36.8 54.0 - 33.6 42.4 57.9 68.3
Stepwise [25] 2017 19.6 41.2 55.5 - 46.7 56.2 70.3 79.2
Race [26] 2018 24.5 43.2 57.1 62.1 - - - -
EUG [23] 2018 42.4 62.6 74.9 - 63.2 72.7 84.1 -
ProLearn [21] 2019 42.6 62.8 75.2 80.4 63.3 72.9 84.3 88.3
SPUE-Net [64] 2023 46.3 66.7 78.1 81.6 65.0 73.3 84.8 89.0
DAL [28] 2018 Cameras 23.0 49.3 65.9 72.2 - - - -
UTAL [27] 2019 35.2 49.9 66.4 - - - - -
uPMnet [19] 2022 - - - - 76.9 83.6 93.1 -
UAAL [22] 2022 60.1 73.2 86.3 89.5 87.0 89.7 97.0 98.6
OIM [65] 2017 None 13.5 33.7 48.1 54.8 43.8 51.1 70.5 76.2
TSSL [36] 2020 30.5 56.3 - - 64.6 73.9 - -
DBC [38] 2019 31.7 58.5 70.1 73.5 67.4 75.6 88.5 91.0
BUC [35] 2019 38.0 61.1 75.1 80.0 61.9 69.2 81.1 85.8
NHAC [37] 2021 40.1 61.8 75.3 79.9 76.0 82.8 92.7 95.6
SSL [39] 2020 43.6 62.8 77.2 80.1 69.3 76.4 88.7 91.0
RPE [42] 2023 40.4 63.3 75.4 80.6 71.5 77.8 89.3 91.7
SRC [41] 2022 40.5 62.7 76.1 80.0 76.5 83.0 93.3 95.0
NPSSL [66] 2025 - - - - 88.9 88.7 98.1 -
DBMPCL [67] 2025 71.6 82.0 91.1 93.4 92.7 94.2 98.7 99.4
SSR-C (Ours) 2025 None 78.7 85.6 93.0 95.1 95.3 95.9 99.2 99.6
IV-C Implementation Details
We adopt ResNet-50 [68] pre-trained on ImageNet [69] as our backbone model and use GeM pooling [70] to obtain the feature embedding of each frame. During training, we set the batch size to 32 and total training epochs to 150. Each tracklet is sequentially partitioned every $l=32$ frames to obtain sub-tracklets. For each sub-tracklet, we randomly sample $M=8$ frames with a stride of 4 to form a sub-tracklet segment. Images are resized to $256\times 128$ and augmented by horizontal flipping and random erasing [71]. As for the optimizer, AdamW with a weight decay of 0.0005 is adopted to update the parameters. The learning rate is initialized to $3.5\times 10^{-4}$ with a decay factor of 0.1 every 50 epochs. At the beginning of each epoch, we use DBSCAN [51] and Jaccard distance [72] for clustering with the epseps set to 0.25 and the minimal number of samples for each cluster set to 2. For hyper-parameters, $\lambda$ and $\delta$ are set to 0.1 and 0.7, respectively. Following [50], we simply set $\tau$, $\alpha$, $\gamma_{1}$, and $\gamma_{2}$ to 0.05, 0.1, 0.5, and 0.25, respectively.
IV-D Comparison with the State-of-the-Art
@@
In Tab. I, we present a comparative analysis of our proposed SSR-C against state-of-the-art methods on two large-scale video-based person Re-ID datasets. Relatively speaking, one-shot learning based and tracklet association based methods usually achieve better performance as shown in Tab. I, attributed to their utilization of a few identity annotations or camera metadata for guidance. However, our proposed SSR-C still shows great superiority over them in performance, without relying on any auxiliary information. For example, on MARS, our proposed method significantly outperforms the state-of-the-art one-shot learning based method SPUE-Net [64], which applies a cooperative learning of local uncertainty estimation combined with determinacy estimation, and improves mAP by 32.4%. SSR-C also shows great superiority to the advanced tracklet association based method UAAL [22], which uses the association ranking method to discover positive and negative tracklets pairs for triplet loss training, with an improvement of 18.6% in mAP.
Our proposed SSR-C surpasses all state-of-the-art unsupervised methods, which overlook not only the unexpected noises and variations within tracklets but also abundant useful identity information in tracklets, by a large margin without whistles and bells. For instance, SSR-C surpasses SRC [41], which resamples and re-weights the hard frames in tracklets to improve clustering ability, with 18.8%/12.9% mAP/Rank-1 on Duke-V, and surpasses SSL [39], which utilizes re-assigned soft label distribution to learn from similar images with smooth constraint, with 35.1%/22.8% mAP/Rank-1 on MARS.
It is worth noting that, without expensive annotations or cumbersome modules, our method is competitive with advanced supervised methods [17, 20, 16, 62, 63] on both datasets. Recently, NPSSL [66] devises cluster-level and sample-level filters to jointly exploit pseudo labels, while DBMPCL [67] explores the discriminant features of diversity distribution in both frame-level tracklet modeling and tracklet-level discrepant contrastive learning. Differently, we perform refined clustering and model optimization at the sub-tracklet level, leveraging the property of local similarity between tracklets. The significant performance gains of SSR-C demonstrate its effectiveness in generating reliable pseudo labels and enabling more discriminative feature learning.
@@
TABLE II: Ablation studies on the MARS and Duke-V datasets. “SSR” denotes our self-supervised refined clustering, which first generates reliable sub-clusters and then merges them self-supervised. “R./D.R.” and “PM” refer to merging reachable/directly reachable sub-clusters throughout training and the progressive sub-cluster merging strategy, respectively.
@@
TABLE III: Comparison with image-based methods on the MARS and Duke-V datasets. All methods are replicated using DBSCAN for clustering. “Cameras” denotes requiring camera labels. Our proposed tracklet partitioning is adopted for comparison methods on Duke-V to ensure convergence.
Methods Year Extra Info. MARS Duke-V
mAP Rank-1 mAP Rank-1
ICE [32] 2021 Cameras 72.3 81.0 86.6 88.2
IIDCL [73] 2024 61.2 74.2 88.0 88.9
SPCL [30] 2020 None 44.7 59.5 77.0 79.5
PPLR [31] 2022 53.6 65.9 86.1 86.2
Cluster Contrast [40] 2022 54.4 67.8 85.5 86.5
SSR-C (Ours) 2025 None 78.7 85.6 95.3 95.9
IV-D1 More Comparison with Image-based Methods
In Tab. III, we further compare with state-of-the-art image-based unsupervised person Re-ID methods by using DBSCAN for clustering and replicating their results on the video-based MARS and Duke-V datasets. Note that we adopt tracklet partitioning for all these image-based methods on Duke-V, otherwise, they struggle to converge as observed in Tab. II (see detailed discussion in Sec. IV-E). SPCL [30] improves clustering via hybrid memory and reliability criterion, PPLR [31] reduces the pseudo label noise by employing the complementary relationship between global and part features, and Cluster Contrast [40] computes contrastive loss at the cluster level. Differently, our proposed SSR-C leverages self-supervised refined clustering to generate reliable pseudo labels and performs CSC loss for optimization at the sub-tracklet level. SSR-C surpasses them significantly on MARS and Duke-V without whistles and bells.
Moreover, SSR-C surpasses methods that utilizes additional camera information [32, 73, 44] by a large margin on both benchmarks. Among them, ICE [32] leverages inter-instance pairwise similarity scores and camera information to boost contrastive learning, and IIDCL [73] employs camera proxies to discern the disparities and similarities among individuals across inter-camera views. Without relying on extra information, SSR-C outperforms the state-of-the-art image-based method IIDCL [73] by 17.5%/11.4% mAP/Rank-1 on MARS and 7.3%/7.0% mAP/Rank-1 on Duke-V, respectively. Our proposed SSR-C leverages the locality of the noises within a tracklet and the similarity between tracklets, effectively addressing the challenges posed by noises and unexpected variations within tracklets.
It is also worth mentioning that since we do not modify the backbone model or introduce any extra burden modules, our proposed SSR-C has no extra cost during inference compared with the baseline method. SSR-C is efficient in real-world application. During training, when processing all the 8,298 tracklets on MARS, our proposed NFTP module and progressive sub-cluster merging strategy only take less than 9 seconds on one NVIDIA GeForce GTX 1080 Ti.
IV-E Ablation Studies
@@
In this section, we first investigate the effectiveness of each component in our proposed SSR-C framework through detailed ablation experiments in Tabs. II, IV, V and VI. Then, to help understand our proposed method and demonstrate its superiority, we conduct in-depth analyses of different values of hyper-parameters and their impacts in Figs. 4 and 5, and compare the clustering statistics between the baseline and our method in Fig. 6.
IV-E1 Effectiveness of NFTP
@@
Taking the noises and unexpected variations within tracklets into account, our proposed NFTP module enables reliable clustering. As shown in Tab. II, Method 2 with NFTP brings 10.4% mAP improvement over Baseline (Method 1) on MARS, and a more remarkable performance improvement is observed on Duke-V. However, directly applying InfoNCE loss based on the original tracklets fails to converge. We argue that tracklets in Duke-V are much longer than MARS (167.6 v.s. 59.5 frames per tracklet on average), which implies more severe variations in tracklets. If we simply average frames within each tracklet without NFTP, the tracklet features with bias would result in inaccurate pseudo labels, and further cause the degradation of the optimization. In contrast, partitioning each tracklet yields sub-tracklets with relatively mild feature changes, which facilitates the feature-based clustering algorithm to obtain accurate pseudo labels. Given that applying NFTP may also increase the amount of training data, we further investigate its influence in Tab. V.
IV-E2 Effectiveness of CSC Loss and Self-Supervised Refined Clustering
@@
Based on the abundant pseudo labels generated with our self-supervised refined clustering, we design CSC loss for effective optimization with dynamically mined positive samples. As shown in the shaded area in Tab. II, compared with Method 2, incorporating either R. or D.R. sub-cluster merging strategy and training with InfoNCE loss show inferior results on MARS. The reason is that although our proposed self-supervised refined clustering produces reliable optimization guidance, InfoNCE still optimizes the model in the coarse tracklet granularity after violent sub-cluster merging. Besides, fusing sub-clusters can inevitably involve negative identities, especially when the capability of the model is weak at the early training stage. However, as the results of Methods 5 and 6 show, CSC loss can take full advantage of the abundant supervision $\hat{y}$ and $\tilde{y}$ from self-supervised refined clustering. The two complement and promote each other, enhancing the discriminative ability of the model significantly.
IV-E3 Effectiveness of Progressive Sub-Cluster Merging Strategy
@@
Compared with Methods 5 and 6, Ours (Method 7) yields an additional improvement on both datasets. The results demonstrate the effectiveness of our progressive sub-cluster merging strategy in adapting the optimization to the feature extraction capability of the model at different training stages.
@@
TABLE IV: Influence of noise filtering (NF) and tracklet partitioning (TP) in the NFTP module.
@@
Baseline 63.9 75.4 3.2 4.0
Baseline ++ NF 71.4 80.7 7.3 9.5
Baseline ++ TP 72.5 79.8 87.4 89.0
Baseline ++ NFTP 74.3 82.2 91.7 93.6
SSR-C - NF 73.8 81.9 93.8 94.6
@@
IV-E4 Ablation Studies of Noise Filtering and Tracklet Partitioning
@@
We provide more detailed ablation results of our proposed NFTP module, which contains two key designs of noise filtering (NF) and tracklet partitioning (TP). (1) NF is to handle harmful frames as depicted in Fig. 1 (a). In unsupervised learning, noisy frames in tracklets can lead to inaccurate clustering and wrong pseudo labels, thereby hurting model learning. More intuitively, adopting noisy filtering to the baseline in Tab. IV shows its effectiveness. (2) TP is designed to alleviate unexpected variations within tracklets as depicted in Fig. 1 (b). Moreover, it helps explore locality within tracklets and enables our progressive self-supervised refined clustering for reliable pseudo labels. We only adopt tracklet partitioning to the baseline and SSR-C in Tab. IV, and the results show its non-negligible necessity. (3) To promote learning and avoid the potential loss of complementary features within tracklets, we propose a progressive sub-cluster merging strategy to merge sub-tracklets of the same identity effectively. We have shown its effectiveness in Tab. II.
TABLE V: The impact of the different numbers of sampled frames per tracklet (#Frames).
Methods #Frames MARS Duke-V
mAP Rank-1 mAP Rank-1
Baseline 8 63.9 75.4 3.2 4.0
16 40.8 55.7 8.5 11.5
Baseline ++ TP 4 71.8 79.7 87.2 88.5
Baseline ++ NFTP 4 74.2 82.2 92.4 93.7
SSR-C 4 75.5 83.9 93.5 94.3
8 78.7 85.6 95.3 95.9
16 70.9 80.3 94.9 95.3
IV-E5 Influence of the Amount of Sampled Frames
@@
In Tabs. II and IV, we notice a remarkable performance improvement on the Duke-V dataset after applying NFTP (especially its tracklet partitioning design), while the baseline based on the original tracklets fails to converge. We further investigate the influence of the number of sampled frames and NFTP on the Duke-V dataset in Tab. V. Firstly, we try different numbers of sampled frames for each tracklet using the baseline method without applying NFTP. The results show that simply increasing the number of sampled frames for each tracklet is useless for enabling model convergence. However, when incorporating tracklet partition, which alleviates the negative impact of unexpected variations within tracklets, a huge performance improvement is observed. Note that on Duke-V, sampling 16 frames per tracklet (Line 2) is roughly equal to the data volume of sampling 4 frames per tracklet after partition (Line 3), but the performance of the former is still unsatisfactory. The results emphasize the indispensable role of the partition design, which alleviates the tracklet feature bias by dividing the tracklets into sub-tracklets. Additionally, the NFTP module better promotes learning with noisy frames being filtered, as shown in Line 4.
We also investigate the influence of the number of sampled frames on the MARS dataset, where the intra-tracklet variations are relatively small. For the baseline method, setting the number of sampled frames per tracklet large will lead to degraded performance. Excessive sampling would introduce redundant frames for short tracklets, and unexpected variations and noises within tracklets for long tracklets. When introducing TP, we observe a notable improvement in performance. TP helps alleviate the bias in tracklet-level feature representation, especially when different segments of a tracklet contain varying levels of useful information. Further improvements are achieved by incorporating the noise filtering design, which refines the sub-tracklet representations.
We further explore the influence of the different numbers of sampled frames after incorporating NFTP in our SSR-C framework in Tab. V. Note that after applying NFTP, we sample frames in the sub-tracklet level, rather than the original tracklet. Sampling 8 frames per sub-tracklet yields the best performance. This configuration strikes a good balance between data diversity and training efficiency. The results show the significant performance gains of our NFTP design, demonstrating its effectiveness. By adaptively aggregating noise-filtered sub-tracklets, our proposed method unleashes the potential of tracklets and promotes unsupervised video person re-identification effectively. Our proposed method does not have any assumption on the dataset property. It achieves good results on both MARS and Duke-V, although the two datasets have different tracklet distributions.
IV-E6 Influence of Different Pooling Strategies
GeM pooling generalizes max and average pooling and is more robust to noises in the unsupervised scenario, so it is adopted in SSR-C to obtain the feature embedding of each frame. We compare different pooling strategies in Tab. VI and find GeM more effective. Note that by adopting either GAP or GMP, our method still surpasses state-of-the-art methods on both datasets by a large margin, showing its effectiveness and robustness.